There is a notice in a lift saying that the maximum weight allowed is 1200 kg.The weights of passengers are normally distributed with mean of 76 kg and a standard deviation of 9.5 kg 15 passenger get into the lift. (a) Calculate the mean and standard deviation of the total weights of 15 passengers.


1
Expert's answer
2021-06-25T07:07:24-0400

"\mu=76,\sigma=9.5"


Maximum weight allowed=1200 kg


(a) n=15,


Sample mean "\bar{x}=\mu=76kg"


Sample standard deviation "s=\dfrac{\sigma}{\sqrt{n}}=\dfrac{9.5}{\sqrt{15}}=2.452 kg"